Trace the curve:
y2(x-a)=x2(2a-x)
1
Expert's answer
2019-03-29T11:04:52-0400

The given curve can be written as


"y^2={x^2(2a-x) \over (x-a)}"

1.Symmetry

The curve is not symmetrical about the y-axis.

The curve is symmetrical about the x-axis.


"(-y)^2={x^2(2a-x) \over (x-a)}\Longrightarrow (-y)^2=y^2"

The curve is not symmetrical in opposite quadrants.

The curve is not symmetrical about the line y=x.

2. Origin.

The curve passes through the origin


"x=0 \Longrightarrow y=0"

3. Intersection with the coordinate axes.


"y-intercept: x=0\Longrightarrow y=0, point(0, 0)""x-intercept: y=0\Longrightarrow 0={x^2(2a-x) \over (x-a)}""x=0\ or\ x=2a""point(0, 0), point(2a, 0)"

4. First derivative

Take derivative with respect to x of both sides of the equation and use the Chain rule


"{d \over dx}(y^2(x-a))={d \over dx}(x^2(2a-x))"

"2y(x-a){dy \over dx}+y^2=2x(2a-x)-x^2"

"{dy \over dx}={2x(2a-x)-x^2-y^2 \over 2y(2a-x)}"

Thus the tangent at (2a, 0) is parallel to y-axis.

5. Tangents at the origin

The equations of the tangents to the curve at the origin is obtained by equating the lowest degree terms in x and y

in the given equation to zero


"xy^2-ay^2=2ax^2-x^3"

"-ay^2=2ax^2"

"2x^2=-y^2"

No tangent at the origin.


6. Asymptote(s)


"y^2(x-a)=2ax^2-x^3"

Asymptote parallel to y-axis


"Coefficient\ of\ y^2\ is\ (x-a)."

"x-a=0\implies x=a"

x=a is the a symptote parallel to y-axis.

There is no horizontal asymptote.

There is no slant (oblique) asymptote.

7. Regions where no part of the curve lies.


"y^2={x^2(2a-x) \over (x-a)}"

"y=\pm \sqrt{x^2(2a-x) \over (x-a)}"

"{x^2(2a-x) \over (x-a)}\geq0 \Longrightarrow a< x \leq 2a"

Note that x=0 is the point on the curve but no branches passes through (0, 0), such a point is called as isolated point.